OpenStudy (anonymous):

Which function below is the inverse of f(x) = The quantity of two x plus three, over five.?

OpenStudy (shamim):

f(x)=(2x+3)/5

OpenStudy (shamim):

is not it?

OpenStudy (shamim):

let y=f(x) x=f^-1(y)

OpenStudy (shamim):

y=(2x+3)/5 x=(5y-3)/2 f^-1(y)=(5y-3)/2 f^-1(x)=(5x-3)/2

OpenStudy (shamim):

did u get it